# Mixed-Reality Digital Twins: Leveraging the Physical and Virtual Worlds for Hybrid Sim2Real Transition of Multi-Agent Reinforcement Learning Policies

### v5

@@ IV-A1 Training and Simulation Parallelization
Fig. 3 depicts the key performance indicators (KPIs) used to analyze the cooperative MARL training without any domain randomization (i.e., NDR). It was observed that the agents took over 600k steps to understand the collective objective of safe intersection traversal. This is marked by a sustainable increase in the cumulative reward (from $\sim$3 to $\sim$8) as well as episode length (from $\sim$470 to $\sim$600 steps). This is also when the policy entropy (i.e., randomness) fluctuated significantly, signifying that the agents were still learning. After this initial exploration, the agents tried reward hacking by choosing to take a longer time to traverse the intersection. This is marked by an increase in the episode length (from $\sim$600 to $\sim$700 steps) between 700k and 750k steps. We anticipate this to be an effect of the last reward term, which continuously rewarded the agents inversely proportional to their distance from the goal. However, this phase was quickly overcome, since the probability of collision or lane boundary violations increased and the resulting reward was comparatively insignificant. By now, the policy entropy was starting to settle but was still fluctuating a little. Towards the end of 1M steps, the policy converged at a stable cumulative reward ($\sim$8) and episode length ($\sim$600 steps), while settling at a policy entropy of $\sim$1.2. For low- and high-domain randomization (i.e., LDR and HDR), the KPIs followed a similar trend but with increasing fluctuations (especially in policy entropy), owing to the randomized parameters.
From a computing perspective, we analyzed the effect of parallelizing the intersection-traversal environment from a single instance (4 agents) up to 25 instances (100 agents). As depicted in Fig. 2(a)-(b) the reduction in training time (up to 76.3%) was quite non-linear, with a saturating point approaching after 15-20 parallel environments.
Figure 3: Training results for cooperative MARL: (a) cumulative reward, (b) episode length, and (c) policy entropy w.r.t. training steps.
Figure 4: Deployment results for cooperative MARL: (a) A1 and A4 avoid collision, (b) A1 finds a gap between A2 and A3 to reach its goal, and (c) A2 and A3 avoid collision, A4 approaches its goal, and A1 is re-spawned.
@@
Figure 6: Training results for competitive MARL: (a) BC loss, (b) GAIL reward, (c) curiosity reward, (d) extrinsic reward, (e) episode length, and (f) policy entropy w.r.t. training steps.
Figure 7: Deployment results for competitive MARL: (a)-(c) denote three frozen snapshots of a block-block-overtake sequence, and (d)-(f) denote three frozen snapshots of a let-pass-and-overtake sequence.
@@
IV-A2 Deployment and Sim2Real Transfer
The trained policies were first deployed and verified in simulation, where we observed interesting emergent behaviors among the agents. The agents strategically slowed down or steered away from each other to avoid collision (refer Fig. 4).
Next, we quantitatively analyzed the policies trained with different grades of domain randomization (i.e., NDR, LDR, and HDR) and benchmarked them against FGM. The design of experiments followed 16 simulation runs and 16 real-world deployments, where the performance was assessed across 3 KPIs, viz. success rate, cumulative reward, and episode duration aggregated across all the agents (since this was a cooperative scenario) as depicted in Fig. 5. For real-world deployments, our experiments cycled across all the agents, such that each agent was physically deployed in the loop with the simulated environment comprising its virtual peers (refer Section III-D for implementation details).
It was observed that FGM was least successful ($<$30%) across simulation as well as real-world experiments. The same fact was reflected across the reward ($<$5 points) as well as duration ($<$500 steps) metrics. NDR-MARL performed better with mean success rates above 45%, which allowed it to cash in over 6 reward points on average. Here, the episode duration was between 400 and 650 steps in most cases, with outliers depicting early collisions. LDR-MARL was the most consistent with success rates reaching as high as 80%, rewards reaching as high as 9 points, and episode durations ranging between 500 and 600 steps in most of the cases. Lastly, it was observed that HDR-MARL performed poorer than other MARL configurations, where it could only achieve up to 40% success rate.
Finally, it is worth mentioning that the closeness between sim and real metrics estimates the sim2real gap, which was least for LDR-MARL (4.12%) followed by NDR-MARL (9.38%), FGM (16.01%), and HDR-MARL (33.85%).
IV-B Competitive Multi-Agent Scenario
IV-B1 Training and Simulation Parallelization
Fig. 6 depicts the KPIs used to analyze the competitive MARL training without any domain randomization (i.e., NDR). It was observed that the agents initially (until $\sim$200k steps) tried aggressive maneuvers, which mostly resulted in collisions. This is marked by the low extrinsic rewards ($<$50) and episode lengths ($<$1400 steps) in this phase. This can also be attributed to the higher BC loss ($>$0.1) as well as lower curiosity ($<$0.8) and GAIL ($<$9) rewards, indicating that the agents had not even started imitating the demonstrations correctly. However, the pre-recorded demonstrations soon (between $\sim$200k and $\sim$800k steps) guided the agents toward completing multiple laps around the race track. This is marked by an exponential reduction in the BC loss (from $\sim$0.1 to $\sim$0.025) and a progressive increase in the extrinsic (from $\sim$50 to $\sim$57), curiosity (from $\sim$0.8 to $\sim$1.1), and GAIL (from $\sim$9 to $\sim$11) rewards as well as the episode length (from $\sim$1400 to $\sim$1600 steps). It was interestingly observed that the red agent (Agent 1) dominated the blue one (Agent 2) till about 500 steps, after which the latter learned the “competitive spirit” and bridged the performance gap. Towards the end of 1M steps, both the policies converged at stable reward values and episode length, while gradually reducing the policy entropy from $>$0.3 to $<$0.05. Here, the non-zero offset in BC loss indicates that the agents did not over-fit the demonstrations; rather, they explored the state space quite well to maximize the extrinsic reward by adopting aggressive “racing” behaviors. The KPIs followed a similar trend for LDR and HDR variations but with higher fluctuations (especially in policy entropy), owing to the randomized parameters.
From a computing perspective, we analyzed the effect of parallelizing the 2-agent adversarial racing family from a single instance (2 agents) up to 10 such families (20 agents) training in parallel, within the same environment. As observed from Fig. 2(c)-(d) the reduction in training time (up to 49%) was less dramatic in this case, with a saturating point approaching after 10$\times$2 parallel agents.
@@ IV-B2 Deployment and Sim2Real Transfer
The trained policies were first deployed and verified in simulation, where we observed interesting adversarial behaviors (e.g., blocking, baiting, overtaking, etc.). These behaviors conveyed that the agents were explicitly competing, while implicitly coordinating to avoid collisions (refer Fig. 7).
Next, we quantitatively analyzed the policies trained with different grades of domain randomization (i.e., NDR, LDR, and HDR) and benchmarked them against FGM. The design of experiments followed 16 simulation runs and 16 real-world deployments, where the performance was assessed across 3 KPIs, viz. win rate, cumulative reward, and episode duration separately for each agent (since this was a competitive scenario) as depicted in Fig. 8. For real-world deployments, our experiments cycled across all the agents, such that each agent was physically deployed in the loop with the simulated environment comprising its virtual peers (refer Section III-D for implementation details).
It was observed that both agents had the lowest average winning rate ($<$25%) with HDR-MARL, although they secured the least mean reward ($<$25 points) with FGM. This highlighted the difference between losing fairly and losing due to collision, which was also corroborated by the outliers in the duration metric. NDR-MARL provided higher ($\sim$30-45%) winning consistency for either agent, but could not surpass that of LDR-MARL ($\sim$35-47%). The same was reflected by the reward metric, wherein LDR-MARL cashed in slightly more reward ($\sim$60-75 points) than NDR-MARL ($\sim$55-70 points). Both performed equally well on the duration metric ($\sim$1400-1700 steps), however, LDR-MARL was slightly more consistent with lower variance.
Finally, the sim2real gap was least for LDR-MARL (2.88%) followed by NDR-MARL (6.88%), HDR-MARL (8.98%), and FGM (13.48%).
@@ V Conclusion
This work identified two pain points in training and deploying MARL systems, and attempted to address them by proposing a scalable and parallelizable digital twin framework. Two representative case studies were formulated to support the claims: a 4-agent collaborative intersection traversal problem and a 2-agent adversarial head-to-head racing problem. The two problems were deliberately formulated with distinct observation spaces and reward functions, but more importantly, also the learning architecture (vanilla MARL vs. demonstration-guided MARL). We analyzed the training metrics in each case and also noted the non-linear effect of agent/environment parallelization on the training time, with a hardware/software-specific point of diminishing return. Finally, we presented a mixed-reality sim2real transfer of the trained policies using a single physical vehicle, which was immersed within the proposed digital twin framework to interact with its virtual peers in a virtual environment.

### v6

@@ IV-A1 Training and Simulation Parallelization
Fig. 4 depicts the key performance indicators (KPIs) used to analyze the cooperative MARL training without any domain randomization (i.e., NDR). It was observed that the agents took over 600k steps to understand the collective objective of safe intersection traversal. This is marked by a sustainable increase in the cumulative reward (from $\sim$3 to $\sim$8) as well as episode length (from $\sim$470 to $\sim$600 steps). This is also when the policy entropy (i.e., randomness) fluctuated significantly, signifying that the agents were still learning. After this initial exploration, the agents tried reward hacking by choosing to take a longer time to traverse the intersection. This is marked by an increase in the episode length (from $\sim$600 to $\sim$700 steps) between 700k and 750k steps. We anticipate this to be an effect of the last reward term, which continuously rewarded the agents inversely proportional to their distance from the goal. However, this phase was quickly overcome, since the probability of collision or lane boundary violations increased and the resulting reward was comparatively insignificant. By now, the policy entropy was starting to settle but was still fluctuating a little. Towards the end of 1M steps, the policy converged at a stable cumulative reward ($\sim$8) and episode length ($\sim$600 steps), while settling at a policy entropy of $\sim$1.2. For LDR and HDR, the KPIs followed a similar trend but with increasing fluctuations. Finally, for FT, we observed an initial performance degradation (owing to the large and sudden domain gap), which eventually recovered and performed well ($>$60% success rate).
Figure 4: Cooperative MARL training: (a) cumulative reward, (b) episode length, and (c) policy entropy w.r.t. training steps.
From a computing perspective, we analyzed the effect of parallelizing the intersection-traversal environment from 1 replica (4 agents) up to 25 replicas (100 agents). We observed that the throughput decreased linearly up to 347.8 ms as the replicas increased and that the computational workload was highest for handling physics, followed by MARL, rendering, and least (negligible) for the API/GUI calls (refer Fig. 2(b)). As shown in Fig. 2(c), reduction in training time (up to 76.3%) was quite non-linear, with a saturating point approaching after 15-20 parallel environments. This resulted in a boost in MARL sample rate from 78.4 Hz for a single replica to 330.7 Hz for 25 parallel environment replicas (refer Fig. 2(d)).
IV-A2 Deployment and Sim2Real Transfer
The trained policies were first deployed and verified in simulation, where we observed interesting emergent behaviors among the agents. The agents strategically slowed down or steered away from each other to avoid collision. Next, we quantitatively analyzed the policies trained with different grades of domain randomization (i.e., NDR, LDR, and HDR) and benchmarked them against FGM, APF, TEB, and FT. The design of experiments followed 16 simulation runs and 16 real-world deployments, where the performance was assessed across 3 KPIs, viz. success rate, cumulative reward, and episode duration aggregated across all the agents (since this was a cooperative scenario) as depicted in Fig. 5. For real-world deployments, our experiments cycled across all the agents, such that each agent was physically deployed in the loop with the simulated environment comprising its virtual peers (refer Section III-D for implementation details).
@@ IV-A2 Deployment and Sim2Real Transfer
It was observed that FGM was least successful ($<$30%), also reflected across the reward ($<$5 points) as well as duration ($<$500 steps) metrics. APF followed with $\sim$40success rate, $<$5 reward points, and $<$600 steps. TEB averaged with $\sim$30success rate $<$8 reward points, and $<$600 steps, but its performance varied greatly since it returned infeasible solution at times. NDR performed better with mean success rates above 45%, which allowed it to cash in over 6 reward points on average. Here, the episode duration was between 400 and 650 steps in most cases, with outliers depicting early collisions. LDR was the most consistent with success rates reaching as high as 80%, rewards as high as 9 points, and episode durations ranging between 500 and 600 steps. HDR performed poorer than other MARL configurations, where it could only achieve up to 40% success rate. Finally, the performance of base LDR degraded when deployed on turf mat ($<$50% success), which was improved upon fine-tuning ($\sim$75success). It is worth mentioning that the closeness between sim and real metrics estimates the sim2real gap, which was least for LDR (4.12%), followed by NDR (9.38%), TEB (13.32%), FGM (16.01%), APF (19.41%), and HDR (33.85%).
IV-B Competitive Multi-Agent Scenario
Figure 6: Competitive MARL training: (a) BC loss, (b) GAIL reward, (c) curiosity reward, (d) extrinsic reward, (e) episode length, and (f) policy entropy w.r.t. training steps.
IV-B1 Training and Simulation Parallelization
Fig. 6 depicts the KPIs used to analyze the competitive MARL training without any domain randomization (i.e., NDR). It was observed that the agents initially (until $\sim$200k steps) tried aggressive maneuvers, which mostly resulted in collisions. This is marked by the low extrinsic rewards ($<$50) and episode lengths ($<$1400 steps) in this phase. This can also be attributed to the higher BC loss ($>$0.1) as well as lower curiosity ($<$0.8) and GAIL ($<$9) rewards, indicating that the agents had not even started imitating the demonstrations correctly. However, the pre-recorded demonstrations soon (between $\sim$200k and $\sim$800k steps) guided the agents toward completing multiple laps around the race track. This is marked by an exponential reduction in the BC loss (from $\sim$0.1 to $\sim$0.025) and a progressive increase in the extrinsic (from $\sim$50 to $\sim$57), curiosity (from $\sim$0.8 to $\sim$1.1), and GAIL (from $\sim$9 to $\sim$11) rewards as well as the episode length (from $\sim$1400 to $\sim$1600 steps). It was interestingly observed that the red agent (Agent 1) dominated the blue one (Agent 2) till about 500 steps, after which the latter learned the “competitive spirit” and bridged the performance gap. Towards the end of 1M steps, both the policies converged at stable reward values and episode length, while gradually reducing the policy entropy from $>$0.3 to $<$0.05. Here, the non-zero offset in BC loss indicates that the agents did not over-fit the demonstrations; rather, they explored the state space quite well to maximize the extrinsic reward by adopting aggressive “racing” behaviors. The KPIs followed a similar trend for LDR and HDR, but with higher fluctuations (especially in policy entropy), owing to the randomized parameters.
@@ IV-B1 Training and Simulation Parallelization
From a computing perspective, we analyzed the effect of parallelizing 2-agent adversarial racing family up to 10 such families (20 agents) training in parallel, within the same environment. We observed that the throughput decreased quite non-linearly up to 274.2 ms with a dropping rate of change as the replicas increased (indicating higher computational efficiency compared to environment parallelization scheme) and that the computational workload was highest for handling physics, followed by MARL, rendering, and least for the API/GUI calls (refer Fig. 2(f)). As observed from Fig. 2(g), reduction in training time (up to 49%) was less dramatic in this case, with a saturating point approaching beyond 10$\times$2 parallel agents. This resulted in a boost in MARL sample rate from 65.9 Hz for a single replica to 120.3 Hz for 10 parallel agent replicas (refer Fig. 2(h)).
@@ IV-B2 Deployment and Sim2Real Transfer
The trained policies were first deployed and verified in simulation, where we observed interesting adversarial behaviors (e.g., blocking, baiting, overtaking, etc.). These behaviors conveyed that the agents were explicitly competing while implicitly coordinating to avoid collisions. Next, we quantitatively analyzed the policies trained with different grades of domain randomization (i.e., NDR, LDR, and HDR) and benchmarked them against FGM, DEA, and PBC. The design of experiments followed 16 simulation runs and 16 real-world deployments, where the performance was assessed across 3 KPIs, viz. win rate, cumulative reward, and episode duration separately for each agent (since this was a competitive scenario) as depicted in Fig. 7. For real-world deployments, our experiments cycled across all the agents, such that each agent was physically deployed in the loop with the simulated environment comprising its virtual peers (refer Section III-D for implementation details).
It was observed that both agents had the lowest average winning rate ($<$25%) with PBC and HDR, although they secured the least mean reward ($<$25 points) with FGM. This highlighted the difference between losing fairly and losing due to collision, which was also corroborated by the outliers in the duration metric. NDR provided higher ($\sim$30-45%) winning consistency for either agent, but could not surpass that of LDR ($\sim$35-47%), which was similar to DEA. The same was reflected by the reward metric, wherein LDR cashed in slightly more reward ($\sim$60-75 points) than NDR ($\sim$55-70 points). Both performed equally well on the duration metric ($\sim$1400-1700 steps), however, LDR was slightly more consistent with lower variance. DEA could collect most reward amongst non-MARL baselines ($\sim$35-45 points) and also performed more consistently. Finally, the sim2real gap was least for LDR (2.88%), followed by NDR (6.88%), HDR (8.98%), DEA (10.82%), PBC (11.43%) and FGM (13.48%).
@@ V Conclusion
This work identified two pain points in training and deploying MARL systems, and attempted to address them by proposing a scalable and parallelizable digital twin framework. Two representative case studies were formulated to support the claims: a 4-agent collaborative intersection traversal problem and a 2-agent adversarial head-to-head racing problem. The two problems were deliberately formulated with distinct observation spaces and reward functions, but more importantly, also the learning architecture (vanilla MARL vs. demonstration-guided MARL). We analyzed the training metrics in each case and also noted the non-linear effect of agent/environment parallelization on the training time, with a hardware/software-specific point of diminishing return. Finally, we presented a mixed-reality sim2real transfer of the policies (for training/testing) using a single physical vehicle, which was immersed within the proposed digital twin framework to interact with its virtual peers in a virtual environment.